Find the domain and range of \sqrt{x} + 9
Write the answer in interval notation

Answer by lwsshak3(11628)

Find the domain and range of \sqrt{x} + 9
Write the answer in interval notation
***
y=√x+9
x≥0
domain: [0,∞)
..
√x=y-9
x=(y-9)^2
Range: all real numbers, (-∞,∞)